OpenStudy (princess_kylee):

The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (mathmale):

What is the slope of the given line? A line parallel to the given line will have the same slope as does the given line. Once you have this slope, use the point-slope formula to find the equation of the line with this slope and which passes thru the given point. \[y-y _{1}=m(x-x _{1})\]

OpenStudy (princess_kylee):

It doesn't give the slope of the line.

Vocaloid (vocaloid):

what mathmale was trying to say is, you need to find the slope of y - 4 = 2/3 (x - 3)

Vocaloid (vocaloid):

any ideas now?

Vocaloid (vocaloid):

*how?

OpenStudy (princess_kylee):

m= 2/3 ??

Vocaloid (vocaloid):

yes! now, we have m = 2/3, x1 = 1 and x2 = -2 now just plug them into the equation mathmale gave you

Vocaloid (vocaloid):

y−y1=m(x−x1)

Vocaloid (vocaloid):

let me know what you get

OpenStudy (mathmale):

kylee? Do the best you can here with the info you have. Share your work, so that either Vocaloid or I could give you feedback on it.

OpenStudy (princess_kylee):

What do I plug in for y and y1?

OpenStudy (princess_kylee):

1 and -2?

OpenStudy (mathmale):

Your new line is to pass thru the point Q(1,-2). Thus:\[x _{1}=1~and~y _{1}=2\]

OpenStudy (mathmale):

Vocaloid: Looks like we have a conflict? Please resolve it. Thx.

OpenStudy (mathmale):

Excuse me. x1=1 and y1=-2.

Vocaloid (vocaloid):

you're right

Vocaloid (vocaloid):

x1=1 and y1=-2.

OpenStudy (mathmale):

;)

OpenStudy (mathmale):

kylee: Please take this info nd finish writing the equation of this new line.

OpenStudy (princess_kylee):

Does y=0 ?

Vocaloid (vocaloid):

y doesn't equal any specific value, just leave y alone and substitute the rest

Vocaloid (vocaloid):

also leave x alone

OpenStudy (mathmale):

Kylee, you are to write an equation for a straight line here. Both x and y are VARIABLES. There is no reason to let y = 0.

OpenStudy (mathmale):

Instead, replace x1 in the general equation with the given value, 1, and replace y1 in the general equation with the given value, -2. Show y our work, please. Your final result MUST include the variables x and y, in a linear equation / formula.

OpenStudy (princess_kylee):

y-2=2/3(x-1) ?

OpenStudy (mathmale):

At the beginning you are given the equation of a straight line. That equation DOES contain the value of the slope of this line. Your new line must have the same slope, because the new line is parallel to the original line. Vocaloid gave you the correct equation to use: \[y-y _{1}=m(x-x _{1}).\]

OpenStudy (mathmale):

On the left side y ou 've written y-2. This is incorrect because y1 = -2, whereas your y-2 would be correct for y1=2, whereas your y1 is -2. Try again.

OpenStudy (princess_kylee):

I am very confused.

OpenStudy (skullpatrol):

y — ( –2 )

OpenStudy (skullpatrol):

a — b = a + ( –b )

OpenStudy (mathmale):

Yes. The left side of the general equation you're using here is \[y-y _{1}\]

OpenStudy (mathmale):

Two examples: 1) Suppose that y1=2. What would the left side of your equation look like? 2) Suppose that y1=-2. What would the left side of your equation look like? Follow skullpatrol's advice.

OpenStudy (princess_kylee):

y+2=2/3(x-1) ?

OpenStudy (skullpatrol):

Good job!

OpenStudy (mathmale):

Yes. If you'd please enclose that "2/3" in parentheses, you'd be done. 2/3(x-1) is ambiguous. Enclosing "2/3" within parentheses would remove that ambiguity.

OpenStudy (skullpatrol):

^

OpenStudy (princess_kylee):

Does that mean that's my answer to original question?

OpenStudy (mathmale):

You propose a solution 3-4 minutes ago. It was OK and would have been better had you enclosed the slope 2/3 within parentheses.

OpenStudy (princess_kylee):

y+2=(2/3)(x-1) ?

OpenStudy (skullpatrol):

That is the correct answer.

OpenStudy (princess_kylee):

Thank you for all your time and effort! Greatly appreciated everyone!

OpenStudy (skullpatrol):

Thanks for trying to learn something new :-)

OpenStudy (mathmale):

You're welcome. Thx for your perseverance.